OpenStudy (anonymous):

homework eight part one i am okay with the rest but dont understand the first bit

OpenStudy (anonymous):

H8P1 IMPULSE The impulse response is a concept often used in electrical engineering, because it fully characterizes the reaction of a linear system as a function of time. As a first step, we will learn the techniques used to calculate the impulse response of a given system. Consider the circuit shown below. Assume that both the capacitor and the inductor have been left in their relaxed states for a very long time. The voltage source is driven by a signal of vin=δ(t), that is a unit impulse at time t=0. R1=10Ω, L=5mH, R2=10kΩ, C=33nF. (a) What is the value of vR1(t) at t=0−? incorrect (b)Derive a differential equation expression for vR1(t), the voltage across the resistor in series with our inductor. What is the value of vR(t) at t=0+? Hint: it may be helpful to know that ∫0+0−δ(t)dt=1 incorrect (c)Assuming an exponential decay, what is the value for vR(t) at t=1ms? incorrect (d)Similarly, derive a differential equation expression for vC(t), the voltage across our capacitor. What is the value of vC(t) at t=0+? incorrect (e)What is the value for vC(t) at t=1ms?

OpenStudy (anonymous):

Well, I can answer H8P1 (a), which is simply "0V". The time is t = 0-s, so JUST before the Impulse will start. So the supplied voltage is zero and therefore all voltages in the network are 0. Still working on Question (b)-(e).

OpenStudy (anonymous):

ok, I got all of them now. I do all calcs with MatLab since I like scripting :) so: R1 = 10; R2 = 10000; C = 33*10^-9; L = 5 * 10^-3; t0 = 0; t0plus = 1; t0min = 0; tms = 0.001; %Q1 v_R1_0min = R1/L * exp(-t0*R1/L) * t0min % Q2 (see: http://en.wikipedia.org/wiki/RL_circuit#Impulse_response ) v_r1_0plus = R1/L * exp(-t0*R1/L) * t0plus % Q3 (see: http://en.wikipedia.org/wiki/RL_circuit#Impulse_response ) v_r1_1ms = R1/L * exp(-tms*R1/L) * t0plus % Q4 (see: http://en.wikipedia.org/wiki/RC_circuit#Impulse_response ) v_c_0plus = 1/(R2*C) * exp(-t0/(R2*C)) * t0plus % Q5 (see: http://en.wikipedia.org/wiki/RC_circuit#Impulse_response ) v_c_0plus = 1/(R2*C) * exp(-tms/(R2*C)) * t0plus Values then are: v_R1_0min = 0 v_r1_0plus = 2000 v_r1_1ms = 270.6706 v_c_0plus = 3.0303e+003 v_c_0plus = 146.3667

OpenStudy (anonymous):

v_c_0plus is rong dude

OpenStudy (anonymous):

That's scientific notation! v_c_0plus = 3.0303e+003 = 3030.3 Also check your values for R1, R2, C, L and the times. But for me it's correct. Only one point missing for H8P3.

OpenStudy (anonymous):

remove "+" sign its enough